find the standard form of each equation by completing the square then identify the conic section of x^2+y^2-4x+6y-12=0
Answer by lwsshak3(11628)

find the standard form of each equation by completing the square then identify the conic section of x^2+y^2-4x+6y-12=0
..
(x^2-4x+4)+(y^2+6y+9)=12+4+9=25
(x-2)^2+(x+3)^2=25
This is a circle with center at (2,-3) with radius=5